# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v3

@@ 5.2 Conditions for strong IGB
Consider a centered Gaussian distributed dataset that undergoes preprocessing through the standardization $\psi\left(\boldsymbol{\xi^{(a)}}\right)=\boldsymbol{\xi^{(a)}}+\boldsymbol{K},$ where $\boldsymbol{K}$ is an offset with the same dimensionality of the input. The data is then processed through a Multi-Layer Perceptron (MLP) with $L$ hidden layers. This MLP employs ReLU nonlinearities and Max-pooling layers with a kernel size of $m$, and the weights are initialized using the Kaiming normal scheme (with zero biases). In the limit of infinite width for hidden layers, the following scenarios are observed: I. Shifting the center of the input distribution by a vector with the norm $|\boldsymbol{K}|$, even when $L=1$ and $m=1$, results in: $\lim_{|\boldsymbol{K}|\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (18) II. With an increase in the kernel size of the pooling layer, even for a single hidden layer ($L=1$), and in the absence of standardization offset ($|\boldsymbol{K}|=0$), one has: $\lim_{m\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (19) III. In the infinite-depth limit, even in the absence of pooling layers ($m=1$) and standardization ($|\boldsymbol{K}|=0$): $\lim_{L\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (20)
@@ 5.3 Experimental results
Beyond the theoretical analysis presented, in App. I we present experiments demonstrating the presence of IGB also in broader settings. In particular, we explore real data, and various architectures including CNNs, ResNets, MLP-mixers, and Vision Transformers. These experiments demonstrate the presence of IGB in settings commonly used in practice. They also highlight that the use of structured data accentuates IGB, combining the influence of model design with that originating from dataset correlations.
We also provide further experiments on the effect of IGB on dynamics with real data (both binary and multi-class) and different architectures (ResNet, MLP-mixer, and Vision Transformers), showing results qualitatively analogous to those shown in Fig. 2.
@@ 6 Discussion
We examined the classification bias in untrained neural networks, uncovering a phenomenon named IGB that arises from architectural choices that break permutation symmetry among hidden and output nodes within the same layer. The conditions that trigger IGB are related to model design, including data preprocessing, network depth, the selection of activation function and pooling. Factors such as max-pooling, offsetted data standardization or network depth can exacerbate IGB to the extent that all dataset elements are assigned to a single class.
While our analysis is based on random i.i.d. data, we anticipate the presence of IGB in real datasets due to correlated data patches, increasing the likelihood of similar classifications. Hence, we expect stronger IGB in datasets other than random data (as shown in App. I.1 with experiments on real data). Moreover, the hypothesis of unstructured data is instrumental in highlighting the effects induced solely by model design.
@@ Appendix A Notation
  • $\operatorname{erf}\left(\cdot\right)$: Error function.
@@
  • $\boldsymbol{\rho}^{(l;m)}$: vector of the $l^{\text{th}}$ hidden layer, $\left(\rho^{(l;m)}_{0},\dots,\rho^{(l;m)}_{\left\lceil\frac{N_{l}-1}{m}\right\rceil}\right)$, after passing through the max-pooling layer with kernel size $m$; $\rho^{(l;m)}_{i}$ indicate the component corresponding to node $i$.
@@ Appendix C Limit distributions
This section is dedicated to reviewing the convergence in distribution of a sequence of r.v. to its asymptotic distribution. We focus on the conditions that guarantee this convergence and the characterization of the resulting asymptotic distribution.
@@ D.2 Derivation of $\boldsymbol{p_{\mu_{c}}\left({x}\right)}$ for some emblematic cases
$\begin{split}&\left\langle{\rho^{(1;1)}}\right\rangle=\left\langle{\rho^{(1;1)}_{j}}\right\rangle=\int_{\mathbb{R}}x\,p_{g^{(1)}_{j}}^{(\chi)}\left({x}\right)dx=\int_{\mathbb{R}}x\left(\frac{1}{2}\delta(x)+\Theta(x)\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)\right)dx=\\ &\int_{0}^{+\infty}x\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)dx=\sigma_{w}^{2}\int_{0}^{+\infty}\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)d\frac{x^{2}}{2\sigma_{w}^{2}}=\frac{\sigma_{w}}{\sqrt{2\pi}}.\end{split}$ (81)
@@
Figure D.1: Comparison of $\rho^{(1;m)}_{l}$ between the linear activation function (upper plot) and ReLU (bottom plot) across different dataset sizes ($D$). Both plots also include the theoretical distributions derived in App. D.2. Observe the distinct behaviors in the two cases: for the linear activation function, the distribution narrows as $D$ increases, while for the ReLU case, it remains stable. This stability in the ReLU case aligns with expectations from Eq. (81) and Eq. (77). These simulations employed the GB dataset and the SHLP model. For more details, refer to App. I.3.
@@ E.1 Derivation of $\boldsymbol{p_{f_{0}}\left({x}\right)}$
Max-Pool peaks: A practical example for computing $\boldsymbol{p_{f_{0}}\left({x}\right)}$
Utilizing the results from App. E.1, we can calculate the theoretical prediction for $p_{f_{0}}\left({x}\right)$, particularly focusing on the probability density at extreme peaks empirically observed with Max-Pooling.
@@
    $f_{0}=1-\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{\Delta_{\mu}(f_{0})}{2\sigma^{2}_{\infty}}\right)\right)\Rightarrow\Delta_{\mu}(f_{0})=2\sigma^{2}_{\infty}\operatorname{erf}^{-1}\left(2f_{0}-1\right).$ (110)
@@
These computations facilitate a precise understanding of the impact of Max-Pooling on $f_{0}$, especially in terms of evaluating probability densities at the extreme peaks.
@@ E.2 Beyond $\boldsymbol{f_{0}}$: relating the confidence in model assignments to IGB
it implies not only the absence of IGB (wherein the DNN’s classifications distribute the dataset evenly across classes), but also that every assignement has an equal probability for each class.
@@
In conclusion, the approach shown in this section paves the way for more accurate analyses. For example, knowing the $\{C_{O^{(c)}}\}$, we could calculate the entropy $H\left(\{C_{O^{(c)}}\}\mid\mathcal{W},\chi\right)$ and reformulate the absence of IGB as a principle of maximum entropy. However, it is important to note that the presence of the normalization factor complicates the derivation of the distribution of $C_{O^{(c)}}$, which can be avoided by using $R_{C}$. This also provides a natural variable to encapsulate the gap between the assignment variables. Finally, although we focused on the binary problem for simplicity, extensions to the multi-class case are possible using approaches similar to those described in App. H, for example by considering in the ratio the two classes with the highest confidence.
@@ F.3 Eliminate/Trigger IGB with a generic activation function
Figure F.1: Contrast between SReLU and ReLU. Shifting ReLU to meet Eq. (127) conditions eliminates IGB. The figure shows the differing regimes through $p_{f_{0}}\left({x}\right)$. Dataset GB and model SHLP were used for these simulations (see App. I.3 for details).
@@
Appendix G Amplification of IGB level
@@ G.1 Effect of data standardization
with $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$. We then consider an architectural setting without max-pooling and with a linear activation function. As seen in App. D.2, this architectural setting does not lead to the emergence of IGB using independent inputs centered at 0. Repeating the same calculation using the preprocessed dataset, $\psi(\chi)$, according to Eq. (137) we can instead show the appearance of IGB. Indeed, starting from Eq. (75) we have
@@
From which it follows that
@@ G.2 Effect of Max-Pooling
Consider a dataset with Gaussian-distributed, i.i.d. components, where $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$, and $\xi_{b}^{(a)}$ represents the $b$-th component of the $a$-th input vector. Inputs are processed through an MLP with a single hidden layer ($L=1$), mapping the input to output according to Eq. (21), Eq. (22), and Eq. (23). The weights are initialized following the Kaiming normal scheme, i.e., $w^{(l)}_{ij}\sim\mathcal{N}\left(0,\frac{\sigma_{w}^{2}}{N_{l-1}}\right)$, with zero bias weights. In a setting with Max-pooling layer and ReLU activation function, an increase in the kernel size, $m$, leads to an asymptotic amplification of the Initial Guessing Bias (IGB). Formally: $\lim_{m\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (151)
@@
In our proof, we focus on the statistical behavior of the maximum value in a growing set of random variables (r.v.s), particularly emphasizing positive event occurrences. This analysis remains applicable if we substitute the ReLU activation function with a linear one, as ReLU doesn’t change the distribution within the positive domain of its support. This implies that the amplification of Initial Guessing Bias (IGB) in models with Max-Pooling is independent of the choice of activation function, manifesting even without any activation function. This finding isn’t surprising since Appendix D.1 demonstrates that the key factor in the emergence of IGB is the statistics of nodes post-pooling layer, particularly $\left\langle{\rho^{(1;m)}}\right\rangle$. Hence, when considering models with pooling layers, it’s crucial to evaluate the joint impact of the activation function and the pooling layer to accurately assess the presence and intensity of IGB.

### v4

@@ 5.2 Conditions for strong IGB
Consider a centered Gaussian distributed dataset that undergoes preprocessing through the standardization $\psi\left(\boldsymbol{\xi^{(a)}}\right)=\boldsymbol{\xi^{(a)}}+\boldsymbol{K},$ where $\boldsymbol{K}$ is an offset with the same dimensionality of the input. The data is then processed through a Multi-Layer Perceptron (MLP) with $L$ hidden layers. This MLP employs ReLU nonlinearities and max pooling layers with a kernel size of $m$, and the weights are initialized using the Kaiming normal scheme (with zero biases). In the limit of infinite width for hidden layers, the following scenarios are observed: I. Shifting the center of the input distribution by a vector with the norm $|\boldsymbol{K}|$, even when $L=1$ and $m=1$, results in: $\lim_{|\boldsymbol{K}|\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (18) II. With an increase in the kernel size of the pooling layer, even for a single hidden layer ($L=1$), and in the absence of standardization offset ($|\boldsymbol{K}|=0$), one has: $\lim_{m\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (19) III. In the infinite-depth limit, even in the absence of pooling layers ($m=1$) and standardization ($|\boldsymbol{K}|=0$): $\lim_{L\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (20)
@@ 5.3 Experimental results
5.3.1 Robustness of IGB in real settings
Beyond the theoretical analysis presented, in App. I we present experiments demonstrating the presence of IGB in broader settings. In particular, we explore real data and various architectures including CNNs, ResNets, MLP-mixers, and Vision Transformers. These experiments demonstrate the presence of IGB in settings commonly used in practice. They also highlight that the use of structured data accentuates IGB, combining the influence of model design with that originating from dataset correlations. The experiments analyze binary as well as multi-class datasets, both of which are covered by our theory. [^6] While we focus in our discussion on the binary case for the sake of clarity, refer to App. H
5.3.2 Effects of IGB on Training Dynamics
We also provide further experiments on the effect of IGB on learning dynamics with real data (both binary and multi-class) and different architectures (ResNet, MLP-mixer, and Vision Transformers), showing results qualitatively analogous to those shown in Fig. 2. We consider the case of balanced datasets, where the presence of IGB translates into a discrepancy between the fractions of class labels and the corresponding fractions of model predictions. This mismatch poses a natural limit on the performance we can reach, as explained in Sec. 3. In the presence of IGB in these settings, it becomes crucial to reabsorb the predictive bias during the dynamics to achieve good performance. It then becomes important to determine the time required by the dynamics to reabsorb the bias. Our experiments show how this reabsorbing time is critically dependent on the level of IGB; with high levels of IGB, convergence can become dramatically slow. It can also be observed how the improvement in performance over the dynamics proceeds consistently with the absorption of Guessing Bias. Figs. 2, I.4, and I.5 show this by tracking both accuracy and the maximum among the class fraction of guesses, used as a measure of IGB. The comparison of these two trends shows that during the dynamics, the level of guessing bias reduces while the performance improves accordingly.
5.3.3 Effects of IGB on Pre-Trained Models
Figure 6: The distribution $p_{f_{0}}\left({x}\right)$ on pre-trained EfficientNetV2 with the binary dataset (C10). The difference lies in the untrained head: a fully connected layer (PT-ENetV2-FC) versus a MLP (PT-ENetV2-MLP). More details in App. I.4. Increasing head depth amplifies IGB, consistently with our analysis.
We show that in the context of transfer learning, IGB can persist in pre-trained models, where only the final layer(s) of the architecture are untrained. Experiments outlined in App. I.3 (see also Fig. 6) show the presence of IGB in these models and its potential amplification, consistent with our analysis. Acknowledging the presence of IGB in this context is relevant, especially in few-shot learning (Wang et al., 2020), where the fine-tuning dynamics do not involve a large number of steps/epochs.
@@ 6 Discussion
We examined the classification bias in untrained neural networks, uncovering a phenomenon named IGB that arises from architectural choices that break permutation symmetry among hidden and output nodes within the same layer. The conditions that trigger IGB are related to model design, including data preprocessing, network depth, the selection of activation function and pooling. Factors such as max pooling, offsetted data standardization or network depth can exacerbate IGB to the extent that all dataset elements are assigned to a single class.
While our analysis is based on random i.i.d. data, we anticipate the presence of IGB in real datasets due to correlated data patches, increasing the likelihood of similar classifications. Hence, we expect stronger IGB in real datasets (as empirically confirmed in App. I.1). Moreover, the hypothesis of unstructured data is instrumental in highlighting the effects induced solely by model design.
@@
Acknowledgements
This work was supported by the Swiss National Foundation, SNF grant # 196902.
@@ Appendix A Notation
  • $\operatorname{erf}\left(\cdot\right)$: Error function. $\operatorname{erf}\left(x\right)\equiv\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$
@@
  • $\boldsymbol{\rho}^{(l;m)}$: vector of the $l^{\text{th}}$ hidden layer, $\left(\rho^{(l;m)}_{0},\dots,\rho^{(l;m)}_{\left\lceil\frac{N_{l}-1}{m}\right\rceil}\right)$, after passing through the max pooling layer with kernel size $m$; $\rho^{(l;m)}_{i}$ indicate the component corresponding to node $i$.
@@ Appendix C Limit distributions
This section reviews the convergence in distribution of a sequence of r.v. to its asymptotic distribution. We focus on the conditions that guarantee this convergence and the characterization of the resulting asymptotic distribution.
@@ D.2 Derivation of $\boldsymbol{p_{\mu_{c}}\left({x}\right)}$ for some emblematic cases
$\begin{split}&\left\langle{\rho^{(1;1)}}\right\rangle=\left\langle{\rho^{(1;1)}_{j}}\right\rangle=\int_{\mathbb{R}}x\,p_{g^{(1)}_{j}}^{(\chi)}\left({x}\right)dx=\int_{\mathbb{R}}x\left(\frac{1}{2}\delta(x)+\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)\right)dx=\\ &\int_{0}^{+\infty}x\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)dx=\sigma_{w}^{2}\int_{0}^{+\infty}\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)d\frac{x^{2}}{2\sigma_{w}^{2}}=\frac{\sigma_{w}}{\sqrt{2\pi}}.\end{split}$ (81)
@@
Figure D.1: Comparison of the distribution $p_{\mu_{c}}\left({x}\right)$ between the linear activation function (upper plot) and ReLU (bottom plot) across different dataset sizes ($D$). Both plots also include the theoretical distributions derived in App. D.2. Observe the distinct behaviors in the two cases: for the linear activation function, the distribution narrows as $D$ increases, while for the ReLU case, it remains stable. This stability in the ReLU case aligns with expectations from Eq. (81) and Eq. (77). These simulations employed the GB dataset and the SHLP model. For more details, refer to App. I.4.
@@ D.3 Derivation of $\boldsymbol{p_{O^{(c)}}^{(\chi)}\left({x}\right)}$ for some emblematic cases
D.4 Non identically distributed classes
For our analysis we consider unstructured data identically distributed among classes. This choice places us in a symmetric setting where the presence of a predictive bias cannot be linked to differences between classes. This setting allows us to isolate the effect of the architecture sidestepping potential additional effects coming from dataset attributes. However, extending the analysis to other settings is possible; to illustrate this we now show that the analysis can be extended to include classes that are not identically distributed. Specifically, we will extend the current setting by introducing class-specific differences in the variance of the Gaussian blobs. We begin by demonstrating that, given an initialization, both classes exhibit a predictive bias towards the same class. Let’s define a generic value for the class $c$ input variance, $\sigma_{\xi^{(c)}}^{2}$. We can then repeat the computation for each class subgroup as done in App. D.1, specifically for an MLP with ReLU activation. We divide the dataset into subgroups based on class membership (each now following different statistics) and repeat the computation for each. In particular, consider Eq. (68) and assume (without loss of generality) that for a given class $c$, we have a prediction bias towards class 0, i.e.
$\mu^{(c)}_{0}-\mu^{(c)}_{1}=\frac{\sigma_{w}\sigma_{\xi^{(c)}}}{\sqrt{2\pi}}(S_{w_{0}}-S_{w_{1}})>0.$ (98)
This expression comes from substituting Eq. (81) with the class-dependent variance. Then, for a generic class $c^{\prime}$, we will have:
$\mu^{(c^{\prime})}_{0}-\mu^{(c^{\prime})}_{1}=\frac{\sigma_{w}\sigma_{\xi^{(c^{\prime})}}}{\sqrt{2\pi}}(S_{w_{0}}-S_{w_{1}})=\frac{\sigma_{\xi^{(c^{\prime})}}}{\sigma_{\xi^{(c)}}}\mu^{(c)}_{0}-\mu^{(c)}_{1}>0$
indicating that the predictions for elements of class $c^{\prime}$ will also be biased towards class 0.
Thus, for any given initialization, the bias is directed towards the same class for each subgroup. Furthermore, we can show that each subgroup is characterized by the same level of bias, irrespective of $\sigma_{\xi^{(c)}}^{2}$.The ratio between the two variances, $\textrm{Var}_{\mathcal{W}}(\mu_{c^{\prime},c})$ and $\textrm{Var}_{\chi}(O^{(c^{\prime},c)})$, quantifies the level of IGB for the subgroup of datapoints belonging to class $c$. Repeating the computation from App. D.1 (particularly from Eq. (62), Eq. (74), and Eq. (81)), we get:
$\textrm{Var}_{\mathcal{W}}(\mu_{c^{\prime},c})=\frac{\sigma_{w}^{4}\sigma_{\xi^{(c)}}^{2}}{2\pi}$
Similarly, for the variance of the output nodes (from Eq. (72), Eq. (93), and Eq. (94)), we get:
$\textrm{Var}_{\chi}(O^{(c^{\prime}.c)})=\frac{\sigma_{w}^{4}\sigma_{\xi^{(c)}}^{2}(\pi-1)}{2\pi}$
The expressions above confirm that the ratio between the two variances is independent of $\sigma_{\xi^{(c)}}^{2}$, for each subgroup $c$.
@@ E.1 Derivation of $\boldsymbol{p_{f_{0}}\left({x}\right)}$
max pool peaks: A practical example for computing $\boldsymbol{p_{f_{0}}\left({x}\right)}$
Utilizing the results from App. E.1, we can calculate the theoretical prediction for $p_{f_{0}}\left({x}\right)$, particularly focusing on the probability density at extreme peaks empirically observed with max pooling.
@@
    $f_{0}=1-\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{\Delta_{\mu}(f_{0})}{2\sigma_{\infty}}\right)\right)\Rightarrow\Delta_{\mu}(f_{0})=2\sigma_{\infty}\operatorname{erf}^{-1}\left(2f_{0}-1\right).$ (111)
@@
These computations facilitate a precise understanding of the impact of max pooling on $f_{0}$, especially in terms of evaluating probability densities at the extreme peaks.
@@ E.2 Beyond $\boldsymbol{f_{0}}$: relating the confidence in model assignments to IGB
it implies not only the absence of IGB (wherein the DNN’s classifications distribute the dataset evenly across classes), but also that every assignement has an equal probability for each class. This leads us to the concept of Strong Absence of IGB, which we define as follows:
@@
In conclusion, the approach shown in this section paves the way for more accurate analyses. For example, knowing the $\{C_{O^{(c)}}\}$, we could calculate the entropy $H\left(\{C_{O^{(c)}}\}\mid\mathcal{W},\chi\right)$ and reformulate the absence of IGB as a principle of maximum entropy. However, it is important to note that the presence of the normalization factor complicates the derivation of the distribution of $C_{O^{(c)}}$, which can be avoided by using $R_{C}$. This also provides a natural variable to encapsulate the gap between the assignment variables. Finally, although we focused on the binary problem for simplicity, extensions to the multi-class case are possible using approaches similar to those described in App. H, for example by considering the ratio of the two classes with the highest confidence.
E.2.1 Effect of the Softmax temperature
Eq. (114) is the standard softmaxed expression of the output with the temperature parameter $T=1$. Acting on the temperature can effectively reduce IGB. By applying a high-temperature parameter, we observe a narrowing distribution of $R_{C}$ around $R_{C}=1$, indicating a reduction in IGB as $T\to\infty$. However, it is crucial to acknowledge the complex effects of temperature on training dynamics; specifically, a small beta value (high temperature) may affect learning stability [Agarwala et al., 2020]. While further investigation is needed to fully understand the impact of this mitigation strategy on the training process.
@@ F.3 Eliminate/Trigger IGB with a generic activation function
Figure F.1: Contrast between SReLU and ReLU. Shifting ReLU to meet Eq. (128) conditions eliminates IGB. The figure shows the differing regimes through $p_{f_{0}}\left({x}\right)$. Dataset GB and model SHLP were used for these simulations (see App. I.4 for details).
@@
Appendix G Amplification of the IGB level
@@ G.1 Effect of data standardization
with $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$. We then consider an architectural setting without max pooling and with a linear activation function. As seen in App. D.2, this architectural setting does not lead to the emergence of IGB using independent inputs centered at 0. Repeating the same calculation using the preprocessed dataset, $\psi(\chi)$, according to Eq. (138) we can instead show the appearance of IGB. Indeed, starting from Eq. (75) we have
@@
From this, it follows that
@@ G.2 Effect of max pooling
Consider a dataset with Gaussian-distributed, i.i.d. components, where $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$, and $\xi_{b}^{(a)}$ represents the $b$-th component of the $a$-th input vector. Inputs are processed through an MLP with a single hidden layer ($L=1$), mapping the input to output according to Eq. (21), Eq. (22), and Eq. (23). The weights are initialized following the Kaiming normal scheme, i.e., $w^{(l)}_{ij}\sim\mathcal{N}\left(0,\frac{\sigma_{w}^{2}}{N_{l-1}}\right)$, with zero bias weights. In a setting with max pooling layer and ReLU activation function, an increase in the kernel size, $m$, leads to an asymptotic amplification of the Initial Guessing Bias (IGB). Formally: $\lim_{m\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (152)
@@
In the proof of Thm. G.2, we focused on the statistical behavior of the maximum value in a growing set of random variables (r.v.s), particularly emphasizing positive event occurrences. This analysis remains applicable if we substitute the ReLU activation function with a linear one, as ReLU does not change the distribution within the positive domain of its support. This implies that the amplification of Initial Guessing Bias (IGB) in models with max pooling is independent of the choice of activation function, manifesting even without any activation function. This finding is not surprising, given that App. D.1 demonstrates that the key factor in the emergence of IGB is the statistics of nodes post-pooling layer, in particular $\left\langle{\rho^{(1;m)}}\right\rangle$. Hence, when considering models with pooling layers, it’s crucial to evaluate the joint impact of the activation function and the pooling layer to accurately assess the presence and intensity of IGB.